OpenStudy (anonymous):

Graph the exponential function: y = 4^x

OpenStudy (anonymous):

take log of both sides

OpenStudy (anonymous):

\[\ln y=\ln 4^{x}\]

OpenStudy (anonymous):

\[\ln y=x \ln 4\]

OpenStudy (anonymous):

in the form y =mx + c

OpenStudy (anonymous):

This is y=4^x graphed using geogebra.

OpenStudy (anonymous):

this format is new @fourme

OpenStudy (anonymous):

I cant open the document

OpenStudy (anonymous):

Here try this one.

OpenStudy (anonymous):

ok that works thanks!